Main INDEX
Monthly INDEX
PREV
NEXT
User name jpchen
Log entry time 16:50:40 on April16,2008
Entry number 232451
This entry is a followup to: 232248
keyword=cryotarget status last night and now
Last night, power to cryotarget went off caused loss of controls to HP heaters and Fans. Dave came fixed the problem and re-cooled down the cryotarget.
Somehow after the two thermal cycles due ESR crash yesterday and this power trip, we could not get to the same conditions as we were before yesterday, when we could run LH2 at 20K, with Fan at 36 Hz with 10g/s ESR flow while still keep LD2 target cold (23.4K, with 36 Hz Fan).
Now with 10 g/s, we could only run LH2 at 20K with 29Hz and LD2 pretty much warm (51K at HX, fan 20Hz). Seems the LH2 loop fan draw a lot more power than before. We need to watch carefully.
FIGURE 1